# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v2

@@ II Foreground, Background, and Pseudomomentum
We distinguish explicitly between foreground and background degrees of freedom. The foreground is the object of interest, here a single electron or a collection of them. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
@@
Crucially, the same discrete translational symmetry applies to the background (see Fig. 1). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
@@
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
@@
so the corresponding Noether generator is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
@@
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
@@
Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety: the electron changes its pseudomomentum, but the lattice cannot the respond or recoil in kind. A pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that both foreground and background have pseudomomentum, and can exchage it freely. Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
@@
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2). The very real recoil of the atom labeled in red in Fig. 2 may very well prove to have excited no vibrational quanta when measured in a number state basis. This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
Two measurable subsystem momenta and the total momentum
Figure 1 shows two variations on a mechanical system. At the top, two smooth blocks interact frictionlessly. At the bottom, both blocks are corrugated, leading to periodic interaction.
@@
Figure 2: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions.
Finite-window symmetry
@@
III Lattice degrees of freedom and zero modes
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
@@ III Lattice degrees of freedom and zero modes
IV Electron-Lattice interaction
IV.1 Foreground and background Hamiltonians
@@
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are instead generated dynamically through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [9].
@@
Density operator and exact factorization
@@
Density–density interaction and pseudomomentum conservation
@@
$H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (12)
following the formalism of Van Hove [30]. No assumption about plane-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
Using Eq. (11), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i{\mathbf{q}}\cdot\hat{\bm{\rho}}}\,\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (13)
$=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
[…]
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers.
Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involvement and lattice background recoil is allowed. It seems this should matter for measurable properties.
V Elastic internal scattering and the Debye-Waller fraction
In this section we show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode of the lattice is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
Density–density coupling and emergence of the Bloch potential
We take the foreground–background interaction in density–density form,
@@ V Elastic internal scattering and the Debye-Waller fraction
with the tagged-electron density operator
@@
For the background we use the exact center-of-mass factorization
$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (16)
where $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}_{j}^{(0)}+\hat{\mathbf{u}}_{j}$.
Separating the background density operator into its equilibrium expectation value (taken in the background center-of-mass frame) and fluctuations,
@@
the static component produces an effective periodic potential acting on the foreground electron,
@@
The corresponding single-particle foreground Hamiltonian becomes
@@
which is the usual Bloch Hamiltonian. This periodic potential is not assumed a priori, but emerges as the static mean-field component of the underlying density–density coupling.
The remaining part of $H_{\rm int}$ couples the foreground electron to background density fluctuations $\delta\rho_{\rm bg}$, but with the crucial difference from the clamped-lattice formulation: the full operator $\rho_{\rm bg}(-\mathbf{q})$ retains the explicit center-of-mass translation factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, which enforces exact momentum bookkeeping in the finite, untethered crystal.
Initial and final states
For a perturbative treatment we consider product states of foreground and background,
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i}(\{\mathbf{u}_{j}\}),\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f}(\{\mathbf{u}_{j}\}),$ (20)
where $\Psi^{\rm bg}_{i,f}$ are phonon number states or superpositions thereof. The electronic states are Bloch eigenstates of the static lattice potential.
Using the factorized form of the interaction, any nonvanishing matrix element $\langle\Psi_{f}|H_{\rm int}|\Psi_{i}\rangle$ implies the pseudomomentum transfers
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom, regardless of phonon number changes.
Elastic internal channel and the Debye-Waller factor
For a harmonic lattice, the internal displacements admit the normal-mode expansion

### v3

@@ III.1 Defining Pseudomomentum
It is important that we clearly distinguish component parts We distinguish foreground and background degrees of freedom. The foreground is the object of interest, here a single electron, or it could be a collection of them, or a defect, or a lattice phonoon. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
@@
The same discrete translational symmetry applies to the background. Translating the entire lattice while holding the electron fixed leaves the Hamiltonian invariant, and therefore assigns a pseudomomentum to the background as well. (see Fig. 2). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
@@
Figure 2: Sliding Blocks. Untethered, two blocks attract one another and slide frictionlessly along their common long axis. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now not independently conserved; if the boxes are sliding over one another,their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
@@
so the corresponding symmetry is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
@@
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. The total background pseudomomentum decomposes into
$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R}+\hat{\mathbf{P}}_{\rm ph}$ (51)
@@
the center-of-mass (zero-mode) contribution and the internal phonon contribution. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
@@
Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety within the prior framework: the electron changes its pseudomomentum, but as is conventionally understood, the lattice cannot respond or recoil in kind. A pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that both foreground and background have pseudomomentum, and can exchange it freely. Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
Historically, there has been a trend to treat electron scattering in a crystal as if it were Compton scattering in free space. There is no ether, no center of mass of a substrate. This is elegant, but also it misses something important. Here, we restore for a periodic lattice a full set of rules that apply in free space, as one would wish. There is a total psuedomomentum conserved even if an electron encounters a defect. However, unlike in free space, in crystals there is an aether, so to speak, a silent partner: the lattice. It is more than its phonons, it has a center of mass.
Figure 3 reflects the revised understanding and diagram including the participation of the lattice. This collision takes place in an “ether.”
Figure 3: (top) Traditional diagram of electron deflection by phonon emission or absorption. (bottom) Momentum transfer in a perfect crystal when the lattice zero mode is retained. The collision occurs on a dynamical lattice: quasielastic phonons may be produced, while the change in electron momentum is balanced jointly by the phonon momentum and the center-of-mass recoil of the lattice. Note: When momentum $\Delta p=\hbar q$ is transferred to the lattice center of mass, the associated de Broglie wavelength is $\lambda=2\pi/q$, identical to that of the scattering process. The large lattice mass suppresses the recoil energy but not the wavelength or the momentum transfer.
@@
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}=\hat{\mathbf{P}}_{R}+\hat{\mathbf{P}}_{\rm int}$, undergo equal and opposite changes, in exact analogy with collisions in free space. In particular, any change in $\hat{\mathbf{p}}$ necessarily induces a compensating change in the global (zero-mode) contribution $\hat{\mathbf{P}}_{R}$, while changes in the internal degrees of freedom are optional This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 4). The very real recoil of the atom labeled in red in Fig. 4 may very well prove to have excited no vibrational quanta when measured in a number state basis. This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
@@
Figure 4: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions. As with Mössbauer scattering, the red atom has gained momentum and with it the whole lattice, but when measured, the lattice may still be in the same vibrational state as before the collision (elastic scattering).
III.2 Finite-window symmetry
@@
IV Lattice degrees of freedom and zero modes
We have already encountered a zero mode in the minimal model of an electron and an atom chain. section II. The mode there is the angular momentum.
Now we venture out more generally, and find the same large phase space becomes available for elastic and quasielastic scattering processes, a phase space that is simply absent if the center of mass is absent in the Hamiltonian.
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with a total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
@@ IV Lattice degrees of freedom and zero modes
V Electron-Lattice interaction
V.1 Foreground and background Hamiltonians
@@
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are generated through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [10].
@@
V.2 Density operator and exact factorization
@@
V.3 Density–density interaction and pseudomomentum conservation
@@
following the formalism of Van Hove [34]. No assumption about Bloch-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel. This equation is the essence of the term: density-density interaction.
For a tagged foreground electron, we write the aperiodic part of the Bloch waves as
@@
Substituting into Eq. (57) gives
$\boxed{H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}}$ (59)
Equation (59) shows that the interaction depends only on the relative coordinate $\hat{\bm{\rho}}=\hat{\mathbf{r}}-\hat{\mathbf{R}}$. The symmetry of the phase factor $e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$ is the central structural feature: pseudomomentum $\mathbf{q}$ is exchanged between foreground and background with change‘s always equal and opposite.
When the electron momentum changes from $k$ to $k^{\prime}$, the lattice background absorbs the compensating momentum
$\Delta P_{\rm bg}=-\hbar(k^{\prime}-k).$ (60)
This momentum is carried entirely by the center-of-mass (zero mode) of the background. Internal lattice excitations describe only relative motion and carry no net mechanical momentum.
The associated energy change separates into internal and recoil contributions,
$\Delta E_{\rm lat}=\Delta E_{\rm int}+\frac{\Delta P_{R}^{\,2}}{2M_{\rm tot}},$ (61)
where $\Delta P_{R}=-\hbar(k^{\prime}-k)$ is the center-of-mass recoil momentum. For macroscopic $M_{\rm tot}$, the recoil energy is negligible, so large momentum transfer can occur with little or no change in internal lattice energy.
Translational invariance implies that the background degrees of freedom contain a rigid (center-of-mass) coordinate $\mathbf{R}$ with conjugate momentum
$\hat{\mathbf{P}}_{\bf R}=-\,i\hbar\frac{\partial}{\partial\mathbf{R}}.$
The remaining background variables describe internal excitations (phonons or distortions) defined relative to the center of mass. Accordingly, the background Hilbert space factorizes as
$\mathcal{H}_{\rm bg}=\mathcal{H}_{R}\otimes\mathcal{H}_{\rm int}.$
All mechanical momentum of the background resides in the zero mode,
$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R},$ (62)
while the internal degrees of freedom carry no net mechanical momentum.
The interaction operator
$e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$
transfers momentum $\hbar\mathbf{q}$ from the electron to the background as a whole. This momentum is carried entirely by the center-of-mass degree of freedom. Lattice recoil is therefore not an additional dynamical channel but a direct consequence of translational symmetry.
A nonvanishing matrix element of $H_{\rm int}$ requires
so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
Crucially, this momentum exchange is independent of phonon occupation. Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero, so the lattice can absorb finite momentum through center-of-mass recoil without any change in its internal state.
VI Elastic internal scattering and the Debye-Waller fraction
We now show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
VI.1 Emergence of the Bloch potential
Separating the background density operator into its equilibrium expectation value (taken in the lattice center-of-mass frame) and fluctuations,
@@ VI.1 Emergence of the Bloch potential
the static component produces an effective periodic potential
@@
The corresponding foreground Hamiltonian becomes
@@
which is the usual Bloch Hamiltonian. In this way, the periodic potential is not imposed a priori, but emerges as the mean-field component of the translationally invariant density-density interaction in the background center-of-mass frame.
The remaining part of $H_{\rm int}$ describes coupling to fluctuations of the background density. Crucially, it retains the explicit center-of-mass translation factor, ensuring that momentum transfer is accounted for by recoil of the background as a whole, rather than being assigned solely to internal phonon modes.
VI.2 Initial and final states
For a perturbative treatment, we consider product states
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i},\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f},$ (67)
where $\Psi_{n\mathbf{k}}$ are extended-zone Bloch eigenstates of the static lattice Hamiltonian and $\Psi^{\rm bg}_{i,f}$ are arbitrary background states in the full lattice Hilbert space (not restricted to phonon number eigenstates).
VI.3 Interaction and matrix element
From Eq. (59), the interaction is
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (68)
where $\hat{\mathbf{R}}$ is the lattice center-of-mass operator and $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ acts on internal background degrees of freedom.
Between product states the interaction matrix element is
$\mathcal{M}_{fi}$ $=\big\langle n^{\prime}\mathbf{k}^{\prime},\Psi_{f}^{\rm bg}\big|H_{\rm int}\big|n\mathbf{k},\Psi_{i}^{\rm bg}\big\rangle$
$=V(\mathbf{q})\,\langle n^{\prime}\mathbf{k}^{\prime}|e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}|n\mathbf{k}\rangle\,\langle\Psi_{f}^{\rm bg}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle.$ (69)
VI.4 Total pseudomomentum conservation
Translational invariance of the full electron-lattice system implies exact conservation of total pseudomomentum,
$\mathbf{k}^{\prime}-\mathbf{k}=-\big(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i}\big),$ (70)
where $\mathbf{K}^{\rm bg}_{i,f}$ are eigenvalues of the total background pseudomomentum operator, including the zero (center-of-mass) mode.
Equation (70) is the only kinematic constraint imposed by translational symmetry. No additional reciprocal lattice vector is required when electronic states are treated in the extended-zone scheme.
In mechanical form this reads
$\Delta\mathbf{p}=-\Delta\mathbf{P},$ (71)
so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
VII Partition of pseudomomentum transfer
We now examine how the transferred pseudomomentum $\hbar(\mathbf{k}^{\prime}-\mathbf{k})$ is stored within the background.
The background matrix element contains two distinct structures,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\qquad\text{(center-of-mass recoil)},$ (72)
and
$\hat{\rho}^{\rm int}_{-\mathbf{q}}\qquad\text{(internal lattice operator)}.$ (73)
The conservation law (70) constrains only the total background pseudomomentum change $\Delta\mathbf{K}^{\rm bg}.$ It places no requirement on how that momentum is distributed between internal modes and the center-of-mass degree of freedom.
VII.1 Conventional phonon bookkeeping
In the clamped-lattice approximation one neglects the recoil operator and expands background states in phonon number eigenstates. In that restricted description one identifies
$\Delta\mathbf{K}^{\rm bg}=\sum_{\mathbf{q}}\Delta n_{\mathbf{q}}\,\mathbf{q},$ (74)
so that electronic momentum change is compensated entirely by phonon creation or annihilation.
Energy conservation then requires
$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}=\epsilon_{n\mathbf{k}}\pm\hbar\omega_{\mathbf{q}\lambda}.$ (75)
This phonon-counting condition is therefore a consequence of the harmonic number-state basis together with suppression of the center-of-mass degree of freedom. It restricts the dynamics to a low-codimension subset of the full many-body phase space, in which momentum and energy must be matched by discrete phonon quanta.
VII.2 Full Hilbert-space description
In the complete electron-lattice Hilbert space, Eq. (70) remains the only exact constraint. The compensating background momentum is carried entirely by the center-of-mass (zero mode) of the lattice. Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
Thus, the association of background momentum transfer with specific phonon occupation changes is not fundamental, but a consequence of discarding the zero mode. Phonons may accompany a scattering event, but they do not provide the required momentum balance. The zero mode is the unique carrier of mechanical momentum.
In the clamped-lattice formulation, an electronic transition $k\to k^{\prime}$ is restricted to those phonon modes whose wavevector and frequency simultaneously satisfy both momentum and energy matching conditions. In low dimensions or at low temperature, such solutions are sparse and may be absent altogether except through Umklapp. When it is said that Umklapp momentum is “absorbed by the lattice,” the role of the zero mode is implicitly reintroduced for the special case of a reciprocal lattice vector $G$, even though the lattice can in fact absorb arbitrary momentum.
By contrast, when the translational zero mode is retained, momentum conservation is enforced by recoil of the background as a whole, while energy conservation is governed independently by the internal spectrum. The internal lattice states form an exponentially dense many-body manifold, and all configurations within an energy window $|E_{f}-E_{i}|\lesssim\Gamma$ contribute. The number of quasi-elastic channels therefore scales with the many-body density of states and becomes macroscopic in the thermodynamic limit. This vast phase space is absent when the zero mode is omitted.
VII.3 Recoil-dominated channel
When the zero recoil mode is retained, the required pseudomomentum transfer is taken up entirely by the background center-of-mass degree of freedom:
$\langle\mathbf{K}^{\prime}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}|\mathbf{K}\rangle=\delta_{\mathbf{K}^{\prime},\,\mathbf{K}-\mathbf{q}},$ (76)
so that total pseudomomentum conservation $\mathbf{k}^{\prime}-\mathbf{k}=-(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i})$ is satisfied without reference to phonon occupation changes.
The associated recoil energy is